# Electronic correlation effects in the response of graphene and MoS2 monolayers to the impact of highly-charged ions

### v1

@@ 4.5 Ion-induced evolution of the density of states
Before we can investigate the electron energy spectra of the different simulations, we first have to discuss how to obtain them from the nonequilibrium simulation data. The main problem in time-diagonal methods, such as the G1–G2 scheme, is that spectral information about correlation effects is lost when using HF-propagators for the retarded and advanced Green function in the GKBA, cf. Eq. (5). In contrast, in two-time approaches, such as real-time NEGF theory [37], one has direct access to the spectra of a system. Then the time-resolved DOS can be calculated from the spectral function according to [28]
@@
One such method is given by Koopmans’ theorem [46], which states that the ionization energies of a quantum mechanical system can be approximated by the eigenvalues of the single-particle Hamiltonian. Of course, in correlated systems, this holds only approximately and we can only expect qualitative results. Koopmans’ theorem is well suited for our situation, since the single-particle HF-Hamiltonian can be directly calculated from the one-particle density matrix. To do so, we diagonalize the single-particle Hamiltonian
@@
Below, we present simulation results for the ultrafast electron dynamics in a MoS2 monolayer during the impact of a 32-fold charged Xe-ion (impact occurs at time $t=0$). The left (right) part of Fig. 1 shows the spectra before (after) the ion impact. The white areas show the entire density of states whereas the filled areas are the occupied states. Note that, given the energy-time uncertainty, instantaneous nonequilibrium spectra have to be treated with care. Each of the curves has been generated by performing a time average over a small window of width $\Delta t=0.1t_{0}$, centered at the time given in the figures. The initial energy of the ion is $E_{\rm ion}$, corresponding to a velocity of 0.4070.407nm/fs and an interaction time with the cluster, $\tau\approx\frac{d_{\rm eff}}{v}\approx 2.2$fs where $d_{\rm eff}\approx 9$Å [44].
@@
Finally, we are in a position to investigate the influence of second-order Born correlations on the DOS, in comparison to a Hartree-Fock calculation. This comparison is shown for the final state, t=5t=5fs, in Fig. 11. Evidently, correlations lead to a merge of the two tight binding bands, as well as to an increased occupation of higher-energy states.
@@ 5 Discussion
In this paper, we have extended the discussion of Ref. [6] about the theoretical description of the interaction of slow highly-charged ions with monolayers of quantum materials. Our simulations are based on a nonequilibrium Green functions description of the electrons in the monolayer within a Hubbard model. The monolayer is coupled to the ion via an embedding scheme that models resonant charge transfer to the ion leading to ultrafast (partial) neutralization. The main improvements compared to Ref. [6] are that we included electron-electron correlations via second-order Born self-energies. For this purpose, we applied the G1–G2 scheme [21] in combination with a time-linear embedding scheme [33].
@@
One of the key results of Ref. [6] was the experimental observation of a dramatic difference (about a factor 7) in secondary electron emission, caused by the impact of a highly-charged ion, from a monolayer of graphene as compared to a MoS2 monolayer. Our theoretical analysis allows us to confirm that this effect is caused by the different structural and electronic properties of the two materials, in particular the smaller lattice spacing and the much higher electron mobility in graphene, as compared to MoS2. The different behavior of the two materials was demonstrated by detailed results for the time and space dependence of the electrostatic potential induced inside of the monolayer, cf. Figs. 8 and 9. An interesting new result is that the reduction of SEE is confined to fast ionization processes that occur within 3fs after the ion impact. Here, only electrons that are created with sufficiently high kinetic energy are able to leave the material. On the other hand, for slower projectiles, where the interaction duration with the material is longer, e.g., Ref. [5], SEE is predicted to be substantially enhanced, also in MoS2.
@@
The confirmation of the theoretical explanation given in Ref. [6] is important since there the analysis was restricted to time-dependent Hartree-Fock simulations, whereas quantum materials, such as MoS2 are characterized by strong electronic correlations. By including correlations on the second-order Born level we cannot only substantiate the previous results but also directly predict the influence which correlations have on ICD-induced secondary electron emission. As is demonstrated in Fig. 4.a), correlations directly affect the speed at which the charge on the inner ring is restored after ion neutralization: while a HF calculation predicts a sign change of the charge on the inner ring around t=4t=4fs, with correlations taken into account, this happens already around t=3t=3fs.
@@
Due to the high computational effort, the present G1–G2 simulations were restricted to cluster sizes $N_{s}\leq 150$. While finite clusters of quantum materials are an interesting subject in their own right, the comparison with the macroscopic monolayers used in the experiment carries finite size effects. While long simulations are possible, the main bottleneck is the storage of the four-dimensional tensor of the two-particle Green function. We expect that, by using the recently developed quantum fluctuations approach [47, 48, 49], it will be possible in the near future to significantly extend the system size and also consider multiple layers of quantum materials and study their interaction with highly-charged ions [5].
@@
The direct link between the SEE yield and the shape of the induced electrostatic potential was demonstrated in Ref. [6] by means of semiclassical molecular dynamics simulations for the ICD-generated electrons. This procedure allows one to reproduce the experimentally observed ratio of SEE yields of SLG and MoS2. At the same time, presently no quantum mechanical results for the states the ICD electrons are “born” into and their relative population are available what has required to invoke additional approximations for the initial conditions of the MD simulations [6]. For this reason the theoretical energy-resolved SEE spectra presently cannot fully explain the measured spectra. To improve on the initial conditions and to take into account the quantum mechanical ICD rates, e.g., [11, 5], remains an important topic for future studies. Another issue of current investigation is the inclusion of nearest neighbor and higher-order interaction terms, e.g. via the PPP model. On the other hand, it will be interesting to extend the simulations to more bands, e.g. [41, 42], to more accurately capture the electronic structure. Improvements along these lines will allow one to increase the accuracy of the current predictions for SEE generation in quantum materials, a topic that is also of interest to low-temperature plasma applications [31]. Our present analysis based on correlated many-body simulations, has clearly shown the strength of our approach: it allows one to systematically study various combinations of quantum materials, as well as initial charge and energy of the projectiles and, ultimately, to control and optimize the secondary electron emission.

### v2

@@ 4.5 Ion-induced evolution of the density of states
Before we can investigate the electron energy spectra of the different simulations, we first have to discuss how to obtain them from the nonequilibrium simulation data. The main problem in time-diagonal methods, such as the G1–G2 scheme, is that spectral information about correlation effects is lost when using HF-propagators for the retarded and advanced Green function in the GKBA, cf. Eq. (6). In contrast, in two-time approaches, such as real-time NEGF theory [keldysh64], one has direct access to the spectra of a system. Then the time-resolved DOS can be calculated from the spectral function according to [joost_phd_2022]
@@
One such method is given by Koopmans’ theorem [koopmans_uber_1934], which states that the ionization energies of a quantum mechanical system can be approximated by the eigenvalues of the single-particle Hamiltonian. Of course, in correlated systems, this holds only approximately and we can only expect qualitative results. Koopmans’ theorem is well suited for our situation, since the single-particle HF-Hamiltonian can be directly calculated from the one-particle density matrix. To do so, we diagonalize the single-particle Hamiltonian
@@
Below, we present simulation results for the ultrafast electron dynamics in a MoS2 monolayer during the impact of a 32-fold charged Xe-ion (impact occurs at time $t=0$). The left (right) part of Fig. 1 shows the spectra before (after) the ion impact. The white areas show the entire density of states whereas the filled areas are the occupied states. Note that, given the energy-time uncertainty, instantaneous nonequilibrium spectra have to be treated with care. Each of the curves has been generated by performing a time average over a small window of width $\Delta t=0.1t_{0}$, centered at the time given in the figures. The initial energy of the ion is $E_{\rm ion}$, corresponding to a velocity of 0.4070.407nm/fs and an interaction time with the cluster, $\tau\approx\frac{d_{\rm eff}}{v}\approx 2.2$fs where $d_{\rm eff}\approx 9$Å [schwestka_charge_2020].
@@
Finally, we are in a position to investigate the influence of second-order Born correlations on the DOS, in comparison to a Hartree-Fock calculation. This comparison is shown for the final state, t=5t=5fs, in Fig. 11. Evidently, correlations lead to a merging of the two tight binding bands, as well as to an increased occupation of higher-energy states. Concluding this section, we mention that recently an alternative approach to correlated spectra in G1–G2 simulations was proposed by Reeves and Vlcek [reeves_prl_24] which will be tested in forthcoming work.
@@ 5 Discussion
In this paper, we have extended the discussion of Ref. [niggas_prl_22] about the theoretical description of the interaction of slow highly-charged ions with monolayers of quantum materials. Our simulations are based on a nonequilibrium Green functions description of the electrons in the monolayer within a Hubbard model. The monolayer is coupled to the ion via an embedding scheme that models resonant charge transfer to the ion leading to ultrafast (partial) neutralization. The main improvements compared to Ref. [niggas_prl_22] are that we included electron-electron correlations via second-order Born self-energies. For this purpose, we applied the G1–G2 scheme [schluenzen_prl_20] in combination with a time-linear embedding scheme [balzer_prb_23].
@@
One of the key results of Ref. [niggas_prl_22] was the experimental observation of a dramatic difference (about a factor 7) in secondary electron emission, caused by the impact of a highly-charged ion, from a monolayer of graphene as compared to a MoS2 monolayer. Our theoretical analysis allows us to confirm that this effect is caused by the different structural and electronic properties of the two materials, in particular the smaller lattice spacing and the much higher electron mobility in graphene, as compared to MoS2. The different behavior of the two materials was demonstrated by detailed results for the time and space dependence of the electrostatic potential induced inside of the monolayer, cf. Figs. 8 and 9. An interesting new result is that the reduction of SEE is confined to fast ionization processes that occur within 3fs after the ion impact. Here, only electrons that are created with sufficiently high kinetic energy are able to leave the material. On the other hand, for slower projectiles, where the interaction duration with the material is longer, e.g., Ref. [Niggas_phys-com_21], SEE is predicted to be substantially enhanced, also in MoS2.
@@
The confirmation of the theoretical explanation given in Ref. [niggas_prl_22] is important since there the analysis was restricted to time-dependent Hartree-Fock simulations, whereas quantum materials, such as MoS2 are characterized by strong electronic correlations. By including correlations on the second-order Born level we cannot only substantiate the previous results but also directly predict the influence which correlations have on ICD-induced secondary electron emission. As is demonstrated in Fig. 4.a), correlations directly affect the speed at which the charge on the inner ring is restored after ion neutralization: while a HF calculation predicts a sign change of the charge on the inner ring around t=4t=4fs, with correlations taken into account, this happens already around t=3t=3fs. At the same time, second order Born selfenergies provide only the lowest order correlation corrections. Further improvements require an extension to GW selfenergies in order to verify the influence of stronger correlations and of excitonic effects.
@@
Due to the high computational effort, the present G1–G2 simulations were restricted to cluster sizes $N_{s}\leq 150$. While finite clusters of quantum materials are an interesting subject in their own right, the comparison with the macroscopic monolayers used in the experiment carries finite size effects. While long simulations are possible, the main bottleneck is the storage of the four-dimensional tensor of the two-particle Green function. We expect that, by using the recently developed quantum fluctuations approach [schroedter_cmp_22, schroedter_23, schroedter_cpp_24], it will be possible in the near future to significantly extend the system size and also consider multiple layers of quantum materials and study their interaction with highly-charged ions [Niggas_phys-com_21].
@@
The direct link between the SEE yield and the shape of the induced electrostatic potential was demonstrated in Ref. [niggas_prl_22] by means of semiclassical molecular dynamics simulations for the ICD-generated electrons. This procedure allows one to reproduce the experimentally observed ratio of SEE yields of SLG and MoS2. At the same time, presently no quantum mechanical results for the states the ICD electrons are “born” into and their relative population are available what has required to invoke additional approximations for the initial conditions of the MD simulations [niggas_prl_22]. For this reason the theoretical energy-resolved SEE spectra presently cannot fully explain the measured spectra. To improve on the initial conditions and to take into account the quantum mechanical ICD rates, e.g., [averbukh_prl_04, Niggas_phys-com_21], remains an important topic for future studies. Another issue of current investigation is the inclusion of nearest neighbor and higher-order interaction terms, e.g. via the PPP model. On the other hand, it will be interesting to extend the simulations to more bands, e.g. [shanavas_effective_2015, liu_three-band_2013], to more accurately capture the electronic structure. Improvements along these lines will allow one to increase the accuracy of the current predictions for SEE generation in quantum materials, a topic that is also of interest to low-temperature plasma applications [Bonitz_fcse_19]. Our present analysis based on correlated many-body simulations, has clearly shown the strength of our approach: it allows one to systematically study various combinations of quantum materials, as well as initial charge and energy of the projectiles and, ultimately, to control and optimize the secondary electron emission.